# Gradient-Sign Masking for Task Vector Transport Across Pre-Trained Models

### v1

@@ 5.3 Subset Data Selection
Across datasets and budgets, structured selectors (coreset, herding, $k$-medoids) often provide small but consistent gains over random selection in the few-shot regime. Yet, random selection remains a strong baseline, with performance approaching that of structured methods as $b$ increases, while incurring no memory or computation overheads from embedding or distances pre-computation. Importantly, this shows that our approach remains effective even when the subset $\mathcal{D}_{s}$ is chosen at random, validating its applicability in strict few-shot settings where sophisticated selection strategies are infeasible. Trends are stable across seeds, with variance shrinking as $b$ grows. Finally, structured methods require access to the full target dataset $\mathcal{D}$, which is unrealistic in scenarios such as privacy-constrained or large-scale settings.
5.4 Sensitivity to the Scaling Coefficient
Figure 3: Accuracy at different $\alpha$ values for mean and majority voting sign selections.
We investigate the sensitivity of masked transport to the scaling factor $\alpha\in(0,1]$, providing a proxy for how compatible and robust the transported task vector is with the target backbone. In addition to our proposed majority voting strategy, we consider a baseline where the estimated sign is taken as the sign of the averaged gradients (mean). Results are reported in Fig. 3.
6 Conclusions & Future Works
In this work, we show that the sign structure of gradients provides a powerful and reliable proxy for the ideal directions of the loss landscape of a fine-tuned model. The exceptional performance of our oracle-based approach, which uses the signs of the true task vector, validates this core insight, confirming that effective transfer is possible when the transported task vector is aligned with the target model’s local loss geometry. GradFix successfully approximates this oracle, achieving significant gains and outperforming naive transfer by using only a handful of labeled samples to estimate gradient signs. While our approach is highly effective and robust in low-data regimes, the remaining gap between our method and the oracle highlights clear avenues for future research. Specifically, future work could focus on developing more advanced strategies for estimating gradient signs to better approximate the ideal oracle, exploring different data selection heuristics, or applying this framework to other architectures and transfer learning settings.
Reproducibilty Statement
We provide the codebase in supplementary material to replicate of our results. All hyperparameters used in our experiments are detailed in the appendix. Additionally, we provide complete proofs for all claims made in the paper, ensuring that our results and statements can be independently verified.
Appendix
Appendix A Gradient Sign Estimator Guarantee
We formalize why majority-vote estimation of gradient signs provides a reliable proxy for the true gradient direction in the limited-data regime.
Lemma (Majority-Vote Gradient Sign).
Let $g_{i}:=\nabla_{\theta_{i}}\mathcal{L}_{B}(\theta_{B})$ denote the true gradient of the target loss with respect to parameter $\theta_{B}$, and let $g_{i}^{(n)}$ be i.i.d. per-sample gradients:
where $\varepsilon_{i}^{(n)}$ is symmetric noise around zero. Define
$p_{i}:=\Pr\left[\mathrm{sign}(g_{i}^{(n)})=\mathrm{sign}(g_{i})\right],$ (16)
Then, for all $i$ with $g_{i}\neq 0$, $p_{i}>1/2$.
In particular, the majority-vote estimator:
$\hat{s}_{i}:=\mathrm{sign}\left(\sum_{n=1}^{N}\mathrm{sign}(g_{i}^{(n)})\right)$ (17)
recovers the correct sign with probability, via Hoeffding’s inequality, of:
$\Pr[\hat{s}_{i}=\mathrm{sign}(g_{i})]\geq 1-\exp\left(-2N(p_{i}-1/2)^{2}\right).$ (18)
Proof.
We divide the proof into two parts.
Step 1: Bias of single sample signs.
Define the indicator random variable $X_{n}:=\mathbbm{1}\{\mathrm{sign}(g_{i}^{(n)})=\mathrm{sign}(g_{i})\}\in\{0,1\}$. The success probability of a single sample is $p_{i}=\Pr[X_{n}=1]=\Pr[\mathrm{sign}(g_{i}^{(n)})=\mathrm{sign}(g_{i})]$.
$p_{i}$ $=\Pr[\varepsilon_{i}^{(n)}>-g_{i}]$
$=\Pr[\varepsilon_{i}^{(n)}>0]+\Pr[-g_{i}<\varepsilon_{i}^{(n)}<0]$
$=1/2+\Pr[-g_{i}<\varepsilon_{i}^{(n)}<0]$
Since $\Pr[-g_{i}<\varepsilon_{i}^{(n)}<0]>0$, it follows that $p_{i}>1/2$.
Step 2: Concentration of majority vote.
The majority-vote estimator $\hat{s}_{i}$ succeeds if
$\sum_{n=1}^{N}X_{n}>N/2.$ (19)
$\sum_{n=1}^{N}X_{n}-Np_{i}\leq N/2-Np_{i}=-N(p_{i}-1/2)$ (21)
According to Hoeffding’s inequality (Hoeffding, 1963) for a sum of i.i.d. random variables $X_{n}\in[0,1]$, we have:
$\Pr[\hat{s}_{i}=\mathrm{sign}(g_{i})]=\Pr\left[\sum_{n=1}^{N}X_{n}>N/2\right]=1-\Pr\left[\sum_{n=1}^{N}X_{n}\leq N/2\right]$ (24)
Therefore, we obtain the final bound:
$\Pr[\hat{s}_{i}=\mathrm{sign}(g_{i})]\geq 1-\exp\left(-2N(p_{i}-1/2)^{2}\right)$ (25)
This result formalizes the intuition that, under mild assumptions on per-sample gradient noise, the majority-vote sign over a small batch of samples provides a reliable approximation to the true gradient direction. The probability of correct recovery grows exponentially with both the number of samples $N$ and the signal-to-noise ratio $p_{i}-1/2$. In practice, this guarantees that even very few labeled samples suffice to construct a gradient-sign mask that preserves most of the descent-aligned components of the source task vector.
Appendix B Additional results
B.1 Sign Agreement Analysis
Figure 4: Sign agreement per block for ViT-B/16. The heatmaps show the percentage of sign agreements between the $\bm{m}$$^{*}$ and $\bm{m}$} constructed by computing the signs of the gradients at different $|D_{s}|$ budgets.
To understand the nature of the misalignment between source and target models, we show in Fig. 4 a heatmap of gradient sign agreement. Specifically, we computed the element-wise agreement between the signs of the target task vector ($\tau_{B}$) and the estimated gradients at different data budgets.
This heatmap reveals that agreement is not uniform across different layers, and there is no simple, global sign correlation. This observation provides a direct explanation for the failure of naive task vector transfer. Without a mechanism to correct for this misalignment, simply adding the task vector introduces harmful directions that degrade performance. Our method, in contrast, actively addresses this by using the target model’s gradients to construct a mask, ensuring that only the useful components of the task vector are transferred and aligned with the target loss landscape. Tabs. 5 and 6 demonstrate that $\delta^{A}$ consistently matches or outperforms $\theta_{B}^{opt}$ across both ViT-B/16 and ViT-L/14 models, even at larger $|D_{s}|$ budgets. Moreover, $\delta^{A}$ exhibits substantially lower standard deviation across seeds, confirming its robustness to the random choice of supervision data and its efficiency compared to direct fine-tuning under identical data constraints.
B.2 Hypeparameters selection
We evaluated the optimal task vector application coefficient $\alpha$ using the validation set of each dataset. For $\delta^{*}$, the optimal $\alpha$ is equal to $1$ across all datasets. In Tab. 4, we summarize the coefficients corresponding to the optimal performance of $\delta^{A}$ for each dataset. $\theta_{B}^{\text{opt}}$ are obtained for each dataset by training with the AdamW optimizer (learning rate $1\mathrm{e}{-5}$) on the corresponding dataset $D_{s}$ used to compute the gradient mask $\bm{m}$}, with a single step of gradient descent.
Architecture EUROSAT SVHN GTSRB RESISC45 DTD
ViT-B/16 $0.3$ $0.3$ $0.6$ $0.2$ $0.7$
ViT-L/14 $0.4$ $0.5$ $0.5$ $0.2$ $0.5$
Architecture SNLI MNLI RTE QNLI SCITAIL
T5v1.1 $0.3$ $0.1$ $0.2$ $0.1$ $0.7$
Table 4: Optimal hyperparameters for $\delta^{A}$ on different CLIP and T5 architectures.
B.3 Additional vision results
Table 5: Cross-dataset performance of ViT-B/16 (AA: datacomp xl s13b b90k, $B$: laion2b s32b b82k) models averaged across different seeds initializations
Model $|\mathcal{D}_{s}^{c}|$ EUROSAT RESISC45 GTSRB SVHN DTD AVG
$\theta_{B}$ zero-shot $49.41$ $67.98$ $48.29$ $50.58$ $55.96$ 54.4554.45±7.30
$\theta_{B}$ fine-tune $98.70$ $95.66$ $98.65$ $97.45$ $83.19$ 94.7394.73±5.94
$\theta_{B}$ + $\tau_{A}$ $49.58$ $67.87$ $49.31$ $50.84$ $56.27$ 54.7854.78±7.05
$\theta_{B}$ + $\delta^{*}$ (oracle) $95.06$ $87.06$ $82.92$ $92.04$ $71.44$ 85.7185.71±8.30
$\theta_{B}^{opt}$ 1 56.6156.61±6.06 69.2569.25±0.96 56.0856.08±2.87 61.3261.32±4.09 56.2156.21±0.88 59.8959.89±6.05
$\theta_{B}-\delta^{A}$ 1 61.9461.94±0.43 70.0570.05±0.56 60.8860.88±2.85 71.0771.07±1.82 58.3258.32±0.30 64.4564.45±5.47
$\theta_{B}^{opt}$ 2 59.4959.49±1.43 71.2071.20±1.13 61.7061.70±0.76 62.0162.01±4.40 57.0057.00±0.54 62.2962.29±5.31
$\theta_{B}-\delta^{A}$ 2 65.0765.07±1.10 71.4271.42±0.90 64.3364.33±1.05 70.1970.19±4.55 58.5158.51±0.10 65.9665.96±5.06
$\theta_{B}^{opt}$ 5 61.9961.99±7.29 73.0173.01±0.48 63.0863.08±1.41 67.0367.03±3.93 59.6559.65±0.80 64.9564.95±5.81
$\theta_{B}-\delta^{A}$ 5 66.0566.05±1.21 71.5771.57±0.88 66.6166.61±0.42 73.5973.59±0.82 60.0260.02±0.20 67.5767.57±4.96
$\theta_{B}^{opt}$ 10 59.9859.98±3.77 72.2772.27±1.65 64.5464.54±1.12 67.8567.85±1.02 60.9660.96±0.33 65.1265.12±4.97
$\theta_{B}-\delta^{A}$ 10 66.5966.59±1.83 72.0572.05±0.59 66.0266.02±1.59 74.8274.82±1.22 60.1860.18±0.40 67.9367.93±5.38
$\theta_{B}^{opt}$ 20 60.5960.59±3.94 74.2274.22±0.66 65.5165.51±0.80 67.1967.19±0.65 62.5962.59±0.08 66.0266.02±5.10
$\theta_{B}-\delta^{A}$ 20 67.0567.05±0.41 72.2972.29±0.14 66.4266.42±0.47 74.1174.11±0.59 60.9260.92±0.08 68.1568.15±4.84
$\theta_{B}^{opt}$ 50 58.8058.80±2.45 75.8875.88±0.83 64.9164.91±0.53 67.5867.58±2.91 64.1364.13±0.11 66.2666.26±5.97
$\theta_{B}-\delta^{A}$ 50 66.9466.94±0.46 72.2672.26±0.22 66.1366.13±0.14 74.0774.07±1.52 61.3561.35±0.03 68.1568.15±4.75
Table 6: Cross-dataset performance of ViT-L/14 (AA: datacomp xl s13b b90k, $B$: laion2b s32b b82k) models averaged across different seeds initializations
Model $|\mathcal{D}_{s}^{c}|$ EUROSAT RESISC45 GTSRB SVHN DTD AVG
$\theta_{B}$ zero-shot $62.80$ $73.12$ $56.12$ $37.28$ $63.35$ 58.5358.53±12.35
$\theta_{B}$ fine-tune $98.95$ $97.06$ $99.16$ $97.80$ $83.56$ 95.3195.31±6.13
$\theta_{B}$ + $\tau_{A}$ $62.77$ $73.49$ $56.03$ $39.09$ $63.56$ 58.9958.99±11.80
$\theta_{B}$ + $\delta^{*}$ (oracle) $96.75$ $90.30$ $88.65$ $92.60$ $72.66$ 88.1988.19±8.52
$\theta_{B}-\delta^{A}$ 1 69.6769.67±1.44 76.4576.45±1.33 66.8266.82±0.84 70.1570.15±5.18 65.5065.50±0.74 69.7269.72±4.46
$\theta_{B}-\delta^{A}$ 2 74.1074.10±2.00 76.9776.97±0.68 71.5571.55±2.73 54.3154.31±2.54 66.1066.10±0.59 68.6168.61±8.44
$\theta_{B}-\delta^{A}$ 5 75.5975.59±2.24 76.8276.82±0.48 73.1473.14±1.23 74.4174.41±2.22 66.9566.95±0.69 73.3173.31±3.88
$\theta_{B}-\delta^{A}$ 10 73.7473.74±1.58 77.5977.59±0.57 74.9474.94±0.73 75.8875.88±2.80 67.4167.41±0.48 73.7773.77±3.86
$\theta_{B}-\delta^{A}$ 20 74.8774.87±0.71 78.1678.16±0.33 74.9074.90±0.55 75.7975.79±0.90 67.5567.55±0.24 74.1574.15±3.83
$\theta_{B}-\delta^{A}$ 50 74.7574.75±0.89 78.2778.27±0.18 74.6174.61±0.91 76.7976.79±1.26 67.7767.77±0.01 74.2774.27±3.86
Appendix C Datasets and Supervision Proportions
In this section, we provide detailed information about the datasets used in our experiments and compute, for each one, the supervision proportions corresponding to our subset budgets $|\mathcal{D}_{s}|$. Recall that $|\mathcal{D}_{s}|$ denotes the number of labeled examples per class used to estimate gradient signs. The resulting percentages indicate what fraction of the full training set those few-shot budgets represent.
C.1 Visual Datasets
  • EuroSAT: A dataset based on Sentinel-2 satellite images covering 13 spectral bands, consisting of 27 00027\,000 labeled and geo-referenced samples across 10 classes (Helber et al., 2019).
  • SVHN: A real-world image dataset from Google Street View house numbers, containing 73 25773\,257 labeled digits across 10 classes (0–9) (Netzer et al., 2011).
  • GTSRB: The German Traffic Sign Recognition Benchmark, comprising 39 20939\,209 training images and 12 63012\,630 test images across 43 classes (Stallkamp et al., 2011).
  • RESISC45: A scene classification dataset with 31 50031\,500 RGB images $256\times 256$ from Google Earth, covering 45 scene classes with 700 images per class (Cheng et al., 2017).
  • DTD: The Describable Textures Dataset, consisting of $5640$ images organized into 47 categories inspired by human perception (Cimpoi et al., 2014).
Table 7: Supervision proportions for visual datasets. $|\mathcal{D}_{s}|$ denotes examples per class. Each cell shows the total dataset percentage.
$|\mathcal{D}_{s}|$
@@ C.1 Visual Datasets
C.2 Textual Datasets
  • SNLI: Stanford Natural Language Inference dataset, containing 570 000570\,000 sentence pairs labeled for entailment, contradiction, or neutral (Group et al., 2022).
  • MNLI: Multi-Genre Natural Language Inference dataset, comprising 433 000433\,000 sentence pairs annotated with textual entailment information across various genres (Williams et al., 2018).
  • RTE: Recognizing Textual Entailment dataset, with $2490$ examples for training, 277 for validation, and $3000$ for testing, divided into two classes (Wang et al., 2018).
  • QNLI: Question Natural Language Inference dataset, containing 104 743104\,743 training examples divided into two classes (Wang et al., 2018).
  • SCITAIL: A science entailment dataset built from science question answering, with 23 59623\,596 training examples divided into two classes (Khot et al., 2018).
@@ C.2 Textual Datasets
$|\mathcal{D}_{s}^{c}|$
Dataset # Samples Classes $50$

### v3

@@ C.2 Subset Data Selection
Across datasets and budgets, structured selectors (coreset, herding, $k$-medoids) often provide small but consistent gains over random selection in the few-shot regime. Yet, random selection remains a strong baseline, with performance approaching that of structured methods as $b$ increases, while incurring no memory or computation overhead from embedding or distance pre-computation. Importantly, this shows that our approach remains effective even when the subset $\mathcal{D}_{s}$ is chosen at random, validating its applicability in strict few-shot settings where sophisticated selection strategies are infeasible. Trends are stable across seeds, with variance shrinking as $b$ grows. Finally, structured methods require access to the full target dataset $\mathcal{D}$, which is unrealistic in privacy-constrained or large-scale settings.
C.3 Random Task Vector Transport
Table 7: Ablation study comparing ViT-B/16 models with GradFix ($\delta^{A}$) against a randomized task vector baseline ($\delta^{\text{random}}$). The random vector preserves the mean and standard deviation of the source task vector but lacks its structural information.
Model $|\mathcal{D}_{s}^{c}|$ EUROSAT SVHN GTSRB RESISC45 DTD AVG
$\theta_{B}$ zero-shot - $49.41$ $50.58$ $48.29$ $67.98$ $55.96$ $54.44$
$\theta_{B}$ fine-tune - $98.70$ $97.45$ $98.65$ $95.66$ $83.19$ $94.73$
$\theta_{B}$ + $\tau_{A}$ - $49.58$ $50.84$ $49.31$ $67.87$ $56.27$ $54.77$
$\theta_{B}+\delta^{\star}$ - $95.06$ $92.04$ $82.92$ $87.06$ $71.44$ $85.70$
TransFusionTransFusion - $50.12$ $53.26$ $50.24$ $67.99$ $56.70$ $55.66$
$\theta_{B}^{opt}$ 11 $56.61$ $61.32$ $56.08$ $69.25$ $56.21$ $59.89$
$\theta_{B}+\delta^{\text{random}}$ 11 $49.86$ $51.43$ $48.56$ $68.06$ $56.06$ $54.79$
$\theta_{B}+\delta^{A}$ 11 $61.94$ $71.07$ $60.88$ $70.05$ $58.32$ $64.45$
$\theta_{B}^{opt}$ 22 $59.49$ $62.01$ $61.70$ $71.20$ $57.00$ $62.28$
$\theta_{B}+\delta^{\text{random}}$ 22 $50.07$ $51.45$ $48.64$ $68.14$ $56.05$ $54.87$
$\theta_{B}+\delta^{A}$ 22 $65.07$ $70.19$ $64.33$ $71.42$ $58.51$ $65.90$
$\theta_{B}^{opt}$ 55 $61.99$ $67.03$ $63.08$ $73.01$ $59.65$ $64.95$
$\theta_{B}+\delta^{\text{random}}$ 55 $50.24$ $51.79$ $48.65$ $68.11$ $57.74$ $55.31$
$\theta_{B}+\delta^{A}$ 55 $66.05$ $73.59$ $66.61$ $71.57$ $60.02$ $67.57$
To assess whether the effectiveness of GradFix arises from the transfer of meaningful structural knowledge from the source task vector $\tau_{A}$, rather than solely from the gradient masking mechanism, we performed an ablation using a random task vector. Specifically, we replaced the original source vector $\tau_{A}$ with a randomly generated vector $\tau_{\text{random}}$, drawn from a normal distribution matched to the element-wise statistics of $\tau_{A}$ to ensure a fair comparison:
$\tau_{\text{random}}\sim\mathcal{N}(\mu_{\tau_{A}},\sigma_{\tau_{A}}^{2})$ (27)
where $\mu_{\tau_{A}}$ and $\sigma_{\tau_{A}}$ are the mean and standard deviation of the parameters in $\tau_{A}$. We then applied our proposed gradient-sign masking method to this random vector using the same target gradient mask $\bm{m}$}. The resulting update is defined as:
$\theta_{B}^{\text{rand}}=\theta_{B}+\alpha({\bm{m}}\odot\tau_{\text{random}})$ (28)
We evaluated this baseline using supervision budgets of $|\mathcal{D}_{s}^{c}|\in\{1,2,5\}$ samples per class.
The results are reported in Tab. 7. Notably, the performance of the random vector is often comparable to, or only marginally better than, the zero-shot baseline ($\theta_{B}$). These findings confirm that the gradient-sign mask $\bm{m}$} alone is insufficient to induce strong performance; it must act on a vector that contains valid task-specific information, as provided by $\tau_{A}$.
C.4 Generalization and Robustness
To ensure that merging the transported task vector does not compromise the robust capabilities of the new backbone, we evaluated performance on a held-out support set, ImageNet-R (hendrycks2020many). This allows us to assess whether the target model’s superior zero-shot capabilities are preserved after transfer.
As shown in Tab. 8, transferring the task vector via GradFix ($\theta_{B}+\delta^{A}$) substantially improves accuracy on the downstream tasks compared to the zero-shot baseline, while fully preserving the superior zero-shot generalization of the target model. This confirms that our method effectively adapts the model to the target task without sacrificing the general robustness of the updated pre-training.
Table 8: Generalization analysis. We report accuracy on each downstream task and on the held-out ImageNet-R support set. GradFix improves task performance while preserving the target model’s zero-shot capabilities.
Method EUROSAT SVHN GTSRB RESISC45 DTD
Task Supp. Task Supp. Task Supp. Task Supp. Task Supp.
$\theta_{A}$ zero-shot $49.08$ $68.80$ $47.00$ $68.80$ $43.37$ $68.80$ $58.94$ $68.80$ $47.50$ $68.80$
$\theta_{B}$ zero-shot $49.41$ $79.78$ $50.58$ $79.78$ $48.29$ $79.78$ $67.98$ $79.78$ $55.96$ $79.78$
$\theta_{A}$ ft $98.58$ $65.93$ $93.59$ $67.77$ $98.23$ $65.98$ $92.43$ $65.35$ $79.04$ $65.02$
$\theta_{B}+\delta^{A}$ $67.80$ $79.77$ $69.74$ $79.73$ $65.23$ $80.17$ $72.14$ $79.88$ $59.63$ $79.27$
Appendix D Datasets and Supervision Proportions
In this section, we provide detailed information about the datasets used in our experiments and compute, for each one, the supervision proportions corresponding to our subset budgets $|\mathcal{D}_{s}^{c}|$. Recall that $|\mathcal{D}_{s}^{c}|$ denotes the number of labeled examples per class used to estimate anti-gradient signs. The resulting percentages indicate what fraction of the full training set those few-shot budgets represent.
D.1 Visual Datasets
  • EuroSAT: A dataset based on Sentinel-2 satellite images covering 13 spectral bands, consisting of 27 00027\,000 labeled and geo-referenced samples across 10 classes (eurosat).
  • SVHN: A real-world image dataset from Google Street View house numbers, containing 73 25773\,257 labeled digits across 10 classes (svhn).
  • GTSRB: The German Traffic Sign Recognition Benchmark, comprising 39 20939\,209 training images and 12 63012\,630 test images across 43 classes (gtsrb).
  • RESISC45: A scene classification dataset with 31 50031\,500 RGB images $256\times 256$ from Google Earth, covering 45 scene classes with 700 images per class (cheng_remote_2017).
  • DTD: The Describable Textures Dataset, consisting of $5640$ images organized into 47 categories inspired by human perception (dtd).
Table 9: Supervision proportions for visual datasets. $|\mathcal{D}_{s}^{c}|$ denotes examples per class. Each cell shows the total dataset percentage.
$|\mathcal{D}_{s}^{c}|$
@@ D.1 Visual Datasets
D.2 Textual Datasets
  • SNLI: Stanford Natural Language Inference dataset, containing 570 000570\,000 sentence pairs labeled for entailment, contradiction, or neutral (stanford2022stanford).
  • MNLI: Multi-Genre Natural Language Inference dataset, comprising 433 000433\,000 sentence pairs annotated with textual entailment information across various genres (mnli).
  • RTE: Recognizing Textual Entailment dataset, with $2490$ examples for training, 277 for validation, and $3000$ for testing, divided into two classes (wang2018glue).
  • QNLI: Question Natural Language Inference dataset, containing 104 743104\,743 training examples divided into two classes (wang2018glue).
  • SCITAIL: A science entailment dataset built from science question answering, with 23 59623\,596 training examples divided into two classes (khot2018scitail).
@@ D.2 Textual Datasets
Dataset # Samples Classes $|\mathcal{D}_{s}^{c}|=50$
@@
Appendix E Hyperparameter Selection
We evaluated the optimal task vector application coefficient $\alpha$ using the validation set of each dataset, following standard practice (ilharco2023task; tsv; marczak2025task). For $\delta^{\star}$, the optimal $\alpha$ is equal to $1$ across all datasets. In Tab. 11, we summarize the coefficients corresponding to the optimal performance of $\delta^{A}$ for each dataset. $\theta_{B}^{\text{opt}}$ is obtained for each dataset by training with the AdamW optimizer (learning rate $1\mathrm{e}{-5}$) on the corresponding subset $\mathcal{D}_{s}$ used to compute the gradient mask $\bm{m}$}, with a single gradient-descent step.
Architecture EUROSAT SVHN GTSRB RESISC45 DTD
ViT-B/16 $0.3$ $0.3$ $0.6$ $0.2$ $0.7$
ViT-L/14 $0.4$ $0.5$ $0.5$ $0.2$ $0.5$
Architecture SNLI MNLI RTE QNLI SCITAIL
T5v1.1 $0.3$ $0.1$ $0.2$ $0.1$ $0.7$
Table 11: Optimal hyperparameters for $\delta^{A}$ across CLIP and T5 architectures.
Appendix F Computational Cost Analysis
Following standard scaling-law approximations (kaplan2020scaling), we adopt the commonly used per-parameter FLOP estimates to compare the computational efficiency of our approach against baselines. Let $P$ denote the number of model parameters. We assume the following operation costs:
  • 1 forward pass: $\approx 2P$ FLOPs
  • 1 backward pass: $\approx 4P$ FLOPs
  • Optimizer update (Adam/AdamW): $\approx 10P$ FLOPs per step (accounting for momentum updates, squaring, bias correction, and parameter write-back).
Given this, we derive the costs for GradFix, the few-shot baseline $\theta_{B}^{opt}$, and full fine-tuning.
GradFix cost. GradFix requires a single gradient computation on the target model followed by the masking operation.
  • Forward + Backward pass: $6P$ FLOPs
  • Mask construction and masked task-vector application (element-wise multiplication and addition): $\approx 2P$ FLOPs
Total GradFix Cost: $\approx 8P$ FLOPs.
$\mathbf{\theta_{B}^{opt}}$ cost (one-step fine-tuning). The $\theta_{B}^{opt}$ baseline represents a one-step fine-tuning update on the target model.
  • Forward + Backward pass: $6P$ FLOPs
  • Adam optimizer update: $10P$ FLOPs
Total $\theta_{B}^{opt}$ Cost: $\approx 16P$ FLOPs.
Full fine-tuning cost. Standard fine-tuning (used to produce the original source task vectors) typically involves 2000 training steps.
  • Per step (Forward + Backward + Optimizer): $6P+10P=16P$ FLOPs
  • For 2000 steps: $16P\times 2000$
Comparing the methods reveals significant efficiency gaps:
  • $\theta_{B}^{opt}$ vs. GradFix: $\approx 2\times$ more FLOPs
  • Full fine-tuning vs. GradFix: $\approx 4,000\times$ more FLOPs
  • Full fine-tuning vs. $\theta_{B}^{opt}$: $\approx 2,000\times$ more FLOPs
Euro